Solve the system of equations.
3x + 6y = 3.75
0.375x + 3y = 2.15625

Answer by mangopeeler07(462)

Warning: This solution contains hideous decimals and vicious numbers. Viewer discretion advised.

3x + 6y = 3.75
0.375x + 3y = 2.15625

Multiply the second equation by -2
-0.75-6y=-4.3125

Add .75 to both sides
-6y=-3.5625

Divide by -6
y=.59375

3x + 6y = 3.75

Plug in y
3x+6(.59375)=3.75

Multiply
3x+3.9625=3.75

Subtract 3.9625 from both sides
3x=-0.2125

Divide by 3
x=-0.07083

So,
y=.59375
x=-0.07083